OpenStudy (anonymous):

Please Help! log(base 3)(3^x*y^5/z^1/3) it needs to be simplified

OpenStudy (anonymous):

first use the rule log(A*B) = log(A) + log(B)

OpenStudy (anonymous):

oh damn i forgot all the log crap

OpenStudy (anonymous):

wait shouldnt i use log(a/b)=log(a)-log(b) first?

OpenStudy (anonymous):

it's the same rule really... for the sake of brevity, just consider 1/z^(1/3) as z^(-1/3) ... save you some time and effort

OpenStudy (anonymous):

hold on do i distribute(?) the z^1/3 to both of the things in the numerator?

OpenStudy (anonymous):

i just forgot all that stuff cuz i have been doing a bunch of math classes

OpenStudy (anonymous):

uh sure ... what I think you mean is: "is it ok to just use (3^x * y^5 *z^(-1/3) )" it is ok to just use that like : log (A*B*C) = log(A) + log(B) + log(C)

OpenStudy (anonymous):

oh wait should i bring the z up top?

OpenStudy (anonymous):

ok well i did it like you showed me with a b and c. There is no more simplification i can do with that is there?

OpenStudy (anonymous):

oh wait i remember you can multiply them by the logs of their exponents